can u write me a porportion problem down and check it 2 see if it is write?
Answer by stanbon(75887)

can u write me a porportion problem down and check it 2 see if it is write?
-------
x/8 = 14/100
cross-multiply
x = 8(14/100)
multiply
x = 112/100
reduce
x = 28/25
-----
Checking:
Substitute 28/25 in place of "x" to get:
(28/25)/8 = 14/100
simplify the left side
28/200 = 14/100
reduce both sides
7/50 = 7/50
=======================
Cheers,
Stan H.